Learning Objectives As part of this activity, you want to be able to: Experimentally verify that the strength of the Coulomb force between two charged bodies varies inversely with the square of the separation distance between them. Experimentally verify that the strength of the Coulomb force between two charged bodies varies with the product of the charges. 1.1 Pre-lab Gravitational force Last semester you may have encountered Newtons universal law of gravitation. This law, which describes the force between two point like masses, is mi m2 R2 where, G is Newtons gravitational constant and has an approximate value of 6.674x 10-11 m3/kg s2. mi, m2 are the masses of particles 1 and 2 respectively, R is the distance between the two particles. Although this equation is valid for point-like particles, for R r (r is the radius of the larger of the two objects) this equation gives a good approximation. For the Earth-moon system R is about 60 times the radius of the Earth, so we can expect a reasonable result
1438.10 2 1.01 10 -10 42.24.10-11 1.63.10-11 6 1.17-10-11 7ー -1.60.10-T B 622.10-12 9 5.54 10 10-t-3.19.10-TT. 11-T 3.53-10 12 4.55.10-T3 13 5.18 10-13 14 2.24·10- 15 3.13 10-13 Table 1.1 -Gravitational data 1.1] Question: Use your resources to help you answer these questions. Your mechanics text and your mechanics lab manual may be useful. Be sure to use the distance from the center of the objects. 1. Find the masses of the Earth and the moon and the mean distance between them, then use these values to calculate the force between the Earth and the moon. As you would guess this is a large force due to the relatively close distance between the Earth and the moon and the large mass of the Earth. Would you expect the electric force between the Earth and the moon to be larger or smaller? What is your reasoning? What is a fundamental difference between the mass of an object and the charge on an object that could contribute to these forces being much different? 2. 3. To investigate this model you will fit a set of data to discover the inverse R relation. [1.2] Question: Consider a pair of masses, mi 2kg, m2 3.5kg. An independent researcher performed an experiment to measure the gravitational force between these two masses as a function of distance. Their results are provided in Table 1.1. Using Excel, make a graph of Fg versus R for this data, including an 8% experimental uncertainty in the Y-direction (Y-error bars) I [1.3] Question: Comment on the shape of the graph. Laboratory Manual
1.4] Question: Use Excel to compute the gravitational force between the masses as a function of distance using Eq. 1.1. Plot your results on the same graph as your data. Make sure this model appears as a line. Comment on the fit between your model and your data. Does your data appear to have any systematic error in it? Are you convinced from this experiment that Suppose you werent sure whether Fg o R2. Perhaps it could be F, α R or F R. You could model all three possibilities and compare the results. [1.5] Question: On the same graph, plot two additional lines, based on the following two equations: Based on a comparison of your model lines, which relationship most accu- rately represents your data? In your completed pre-lab, please include only one graph, showing your data with error bars and all three model lines. One final point; the relationship doesnt have to be based on an integer power of [1.6] Question: Try replacing R2 with a few non-integer values, such as R1.5 RLBR22, R5. Based on the quality of your data, can you clearly exclude any of these possibilities? If so, which? Of course, when dealing with non-integer powers, the units become quite odd. 1.2 Introduction In this weeks experiment, you will investigate the force on charged spheres. You will begin by generating a model for the electric force by guessing that the force follows a similar behavior to that of a known force. This new force law can then be tested using the Coulomb balance setup. 1.3 Apparatus High voltage: In this activity, you will use a power supply with a high voltage output in the kilovolt range. Although the current of the power supply is low you can still get a shock from the test leads. Do not touch the metal part of the red lead and do not let the red lead touch anything other than the spheres. Keep the power supply turned off when you are not using it.
Answer #1

[1.1]

1. Mass of Earth (m1)=5.972 × 1024 kg

Mass of Moon(m2)=7.35 x 1022kg

Distance between Earth and Moon(R)= 382,500,000 mts=3.82 × 108 m

G = 6.67 × 10-11 m3/kg.s2

Fg = G m1m2/R2 = 6.67 × 10-11 m3/kg.s2  x 5.972 × 1024 kg x 7.35 x 1022kg / (3.82 × 108 m)2

=2× 1020 N (approx.)

2. The electric force between Earth and moon will be smaller, as electric force takes place between charged body only. As we know that both earth and moon have negligibe charge, so the electric force between them will also be negligible.

3. The fundamental difference between mass and charge of an object is that mass is the measure of inertia of a body where as charge is the measure of electrical property of the body. Also, mass of a body can never be zero, where as the charge can be zero, which can result in such big difference in gravitaional and electric force between two body.